Find the length of a rectangular lot with a perimeter of 132 meters if the length is 4 meters more than the width.
Answer by mananth(16946)

width =w
length = w+4
Perimeter = 2(L+W)
Perimeter = 2(w+w+4)
perimeter = 2(2w+4)
132=4w+8
4w=124
/4
w=31 m the width
length = 35 m